Chapter 4-7 In class assignment- submitted in Canvas

  1. The company’s bank statement shows a cash balance of $15,000. Comparing the company's cash records with the monthly bank statement reveals several additional cash transactions such as checks outstanding of $4,500, deposits outstanding of $5,000, NSF check of $250, and service fee of $60 and a $1,000 note receivable collected by the bank. Calculate the correct balance of cash: $__________

  1. At December 31, the company had account balances in Accounts Receivable of $50,000 and in Allowance for Uncollectible Accounts (AUA) of $600 (credit) before any adjustments. An analysis of the company’s December 31 accounts receivable suggests that the allowance for uncollectible accounts should be 6% of accounts receivable.  After the adjusting entry, what is the balance in the AUA account? $__________

  1. The company had $25,000 of inventory on December 31, Year 2.   During the year they purchased $650,000 and cost of goods was $665,000. What was the company’s beginning inventory for Year 2? $_________

Interest revenue

$2,000

Sales revenue

$120,000

Income tax expense

8,200

Cost of goods sold

62,000

Selling expenses

1,800

Salaries expense

17,500

Utilities expense

900

Rent expense

1,200

  1. Use the following to calculate Net income: $_______
  2. The company purchased new equipment at the beginning of year 1 for $125,000. Management estimated they would use the equipment for 5 years and when they were done using it they estimated the residual value would be $25,000.  What is the equipment’s book valueat the end of year 3 assuming the company uses straight-linedeprecation? $_______
Answer #1

1) Correct Cash balance = 15000+5000-4500 = 15500

2) Balance in AUA Account = 50000*6% = 3000

3) Beginning inventory = 25000+665000-650000 = 40000

4) Calculate net income

Sales revenue 120000
Interest revenue 2000
Total revenue -122000
Expense
cost of goods sold 62000
Salaries expense 17500
Rent expense 1200
Selling expense 1800
Utilities expense 900
Income tax expense 8200
Total expense -91600
Net income 30400

5) Book value =125000-(125000-25000/5*3) = 65000